# Survey on Question Answering over Visually Rich Documents: Methods, Challenges, and Trends

### v1

@@ 6.2 Pretraining Vision-only models
Table 5: Average Normalized Levenshtein Similarity (ANLS) on single and multi-page VQA. ² denotes Single-page-native models concatenating page representations for multi-page; * denotes models using a retriever (PDF-Wukong Xie et al. (2024b) for InternLMXComposer2-4KHD, Naidu et al. (2024) for Pix2Struct, M3DocRAG Cho et al. (2024) for QwenVL and Idefics2). The top-3 scores are in bold.
7 Conclusion and Discussions
LLM-based approaches embedding the VrD representation directly into LLMs through mechanisms, like cross or composite attention improves efficiency, require less training Laurençon et al. (2024) while leveraging LLM capabilities for VrD features (see Section 5).
On the encoder side, we have two possibilities; encoding VrDs with its visual, textual and positional features, or in a vision-only approach. Although LLM-enhanced Vision-only methods (Section 3) are prominent in the recent literature, they face challenges in balancing coarse and fine-grained VrD representations without excessive computational complexity or compression, and are thus not even applicable to multi-page processing without a retriever (see Table 5). We think that approaches leveraging the three modalities (textual, visual and positional) of the document are more efficient, mainly on multi-page use-cases (see Table 5) but remain challenging due to the alignment requirement of those modalities. We believe that studying how models leverage these modalities to answer questions would be valuable to understand the role of each modality in generating responses, whether there are redundancies among them, if one modality dominates the others during generation, and whether they are truly complementary or not.
Based on the results in Table 5, models that maximize the use of positional information appear to achieve the best performance. This suggests that text and layout information alone might be sufficient to answer questions, even for complex charts and figures, provided the layout is optimally taken into account. Leveraging positional biases in attention—especially semantic-based relative positional biases like disentangled attention Peng et al. (2022) or Rich Attention Lee et al. (2022)—proves effective, particularly for tasks like Information Extraction Jaume et al. (2019); Park et al. (2019). Combining positional and semantic alignment of textual and visual modalities seems to allow great performances’ growth Borchmann et al. (2024).
We observed that there is still a challenge of multi-page documents. Retriever-based methods limit inter-page and inter-modality interactions, restricting their ability to perform cross-page reasoning, as we can see on MMLongBenchDoc Ma et al. (2024c) performance as depicted in Table 5. Emerging approaches that encode the document page-by-page with sparse-attention mechanisms for inter-page connections, like global-local or blockwise attention, represent a promising direction. These techniques have demonstrated superior performance in cross-page reasoning tasks Borchmann et al. (2024), positioning them as the future of multi-page document understanding. We also think that slicing approaches that use local information from cropped regions, combined with low-resolution global views Zhang et al. (2024d) could be leveraged for multi-page representation.
Assuming that documents will increasingly be born-digital, we believe the community should focus on models that process documents using text, layout, and optionally visual representation. These models are less costly, more scalable, and particularly more effective for handling multi-page documents.

### v2

@@ 5.3 Pretraining for visual features insertion
Table 3: Average Normalized Levenshtein Similarity (ANLS) on single and multi-page VQA. ² denotes Single-page-native models concatenating page representations for multi-page; * denotes models using a retriever (PDF-Wukong Xie et al. (2024) for InternLMXComposer2-4KHD, Naidu et al. (2024) for Pix2Struct, M3DocRAG Cho et al. (2024) for QwenVL and Idefics2). The top-3 scores are in bold.
6 Conclusion and Discussion
While vision-only methods (Section 3) are gaining prominence in recent literature, they face significant challenges in balancing coarse and fine-grained VrD representations. This often results in excessive computational complexity or compression issues, making these methods unsuitable for multi-page document processing without a retriever (see Table 3). For multi-page understanding, we argue that multi-modal approaches—combining textual, visual, and positional features—are more efficient (see Table 3).
In addition to the computational cost aspect, our view is that the community should prioritize developing methods to handle text, layout, and visual elements in documents, as we observe that documents are increasingly becoming digital-native, with bounding boxes and text readily accessible. However, these approaches remain challenging due to the need for effective alignment across textual and visual features, and due to the need for LLM to handle 2D positional information efficiently.
To reduce redundant information between textual and visual features Tang et al. (2023) and handle both information in an optimal way, we suggest focusing on integrating textual features within the visual representation using cross-attention mechanisms Li et al. (2021b) with text guiding the integration (query) when visual elements are less prominent in the document Borchmann et al. (2024), and visual features guiding when visual elements are major in documents.
Our view is that the community should focus on developing methods to effectively process 2D information, exploring aspects such as granularity, the semantic connection to 2D positions, and multi-level attention mechanisms—both between semantically meaningful blocks and within those blocks, and adapting 2D position encoding to recent approaches Su et al. (2023). As shown in Table 3, models that make extensive use of positional features—such as ERNIE-Layout Peng et al. (2022) and Arctic-TILT Borchmann et al. (2024) – have the best results. This indicates that text and layout information are essential for answering questions, even in complex charts and figures, making efficient layout handling critical.
7 Limitations
A first limitation of our survey lies in the lack of consistent evaluation across different techniques. While we discuss a range of methods—such as 2D position encoding strategies, approaches for integrating visual and textual information, projectors between the visual encoder output and the LLM decoder, sparse attention approaches for multi-page document handling, … – these techniques are evaluated in their original experimental setups, which differ in terms of model architecture, training protocols, and datasets. As a result, it is challenging to draw definitive conclusions about which technique performs best in a given scenario. Although a fairer and more scientifically rigorous comparison would require benchmarking all methods under the same conditions, this was beyond the scope of our survey due to time and resource limitations.
A further limitation of this survey is that most of the comparisons in this survey are based on benchmarks for visual question answering (VQA), while we overlook several traditional document understanding tasks. These tasks include key information extraction, document layout analysis, document classification, or reading order prediction (beyond others), which are essential for many real-world applications such as automatic form processing, contract analysis, and archival document digitization. Our focus on VQA benchmarks is primarily motivated by their widespread use in recent research as a comprehensive testbed for evaluating VrDU approaches both in their information extraction and reasoning capabilities.
Additionally, we focus exclusively on transformer-based approaches. While this choice aligns with the current state of the art, it inevitably excludes earlier yet significant contributions. For instance, traditional methods leveraging LSTMs or Gated Recurrent Units have been widely used in VrDU. More recent work has also started exploring alternative architectures such as state space models Hu et al. (2025). Graph-Based Relationship Modeling approaches, representing documents as hierarchical structures and employing graph neural networks (GNNs) to model relationships between document elements, are also extensively adopted by the community Dai et al. (2024); Zhang et al. (2022); Li et al. (2023b). Due to space and scope constraints, we focused on transformers, which dominate current research and offer a unified framework for integrating visual and textual modalities.
Finally, this survey focuses primarily on generic multi-element documents, such as PDFs and PowerPoint slides, as illustrated in Figure 1 in appendix, rather than specific document types (e.g., tables, charts, or diagrams). Our decision to concentrate on general-purpose documents stems from the desire to provide a broad overview that covers documents combining multiple data types rather than diving into domain-specific challenges. Each specific domain—such as table understanding or chart interpretation—presents its own unique challenges and innovations, like cell, row and columns understanding for table, with approaches modeling column-wise and row-wise self-attention Yin et al. (2020); Deng et al. (2020), derendering tasks for Charts, with approach converting chart image into their Matplotlib code Al-Shetairy et al. (2024) with their associated JSON/CSV Liu et al. (2023), or structure analysis tasks for diagrams, aiming at linking the legend to the diagram content Huang et al. (2024b), which are beyond the scope of this survey.
Appendix A Example Appendix
A.1 Visual Question Answering datasets
We mainly focused on the Visual Question Answering (VQA) task in this survey as a benchmark to compare different models. document-VQA consists of answering a question based on the content of a document, requiring the model to understand both visual and textual information to provide an accurate response.
Datasets type #Pages #Tokens #Tab #Fig Crosspage Unans. Crossdoc #Regions Ans. length
VisualMRC 2021 Wikipedia pages 1.0 151.46 ? ? 9.55
DocVQA 2021b Industry Documents 1.0 182.8 ? ? 2.43
InfographicVQA 2021a Posters (Canva, …) 1.2 217.9 ? ? 1.6
TAT-DQA 2022 Annual Reports 1.3 550.3 >1 ? 3.44
MP-DocVQA 2023 Industry Documents 8.3 2026.6 ? ? 2.2
DUDE 2023 archives, wikimedia 5.7 1831.5 ? ? ✓(2.1%) ✓(12.7%) 3.4
MMLongBenchDoc 2024c ArXiv, Reports, Tuto 47.5 21214.1 25.4% 20.7% ✓(33.0%) ✓(22.5%) 2.8
M3DocVQA 2024 Wikipedia pages 12.2 ? ? ? ? ✓(2.4k) ?
M-LongDoc 2024 Manuals, Reports 210.8 120988 71.8 161.1 180.3
MMDocBench 2024 Multi 1.0 ? ? ? 2.61 4.1
BoundingDocs 2025 Multi 237k ? ? ? >=1 >=1
Table 4: Overview of open-source Question-Answering VrDU datasets on PDFs or PPTs documents, summarizing document characteristics (e.g., average pages, tokens, tabs, figures per document) and question characteristics (e.g., presence of questions requiring cross-pages or cross-documents information, unanswerable questions, and average answer length). #Region refers to the number of regions identified for answering questions in datasets with coordinate annotations. Underlined datasets are standard benchmarks used for model comparison in Table 3.
Figure 1: Illustration of the datasets listed in this survey